OpenStudy (anonymous):

hey...lim as x->1 of ([1/(x-1)]-[1/lnx])...I know it is [(1/0)-(1/0)], what do I do from there?

OpenStudy (anonymous):

Bad way to start. Get a common denominator: [(1*lnx)-(1*(x-1))]/[(x-1)lnx]

OpenStudy (anonymous):

Let a=1/0 a-a=0. Therefore, your answer is 0.

OpenStudy (anonymous):

I got -1/2, anyone agree?

OpenStudy (anonymous):

you got it right. It is -1/2. Using hopitals rule.

OpenStudy (anonymous):

@Harwin: No can do. 1/0 = infinity. Infinity - infinity is one of the 7 indeterminate forms. it's not 0. Esperantist is correct. Get a common denominator. Then it may be possible to use l'Hopital's rule.

OpenStudy (anonymous):

Actually, this is probably the best way-> http://www.wolframalpha.com/input/?i=lim+x-%3E1+%28%28 [1%2F%28x-1%29]-[1%2Flnx]%29%29

OpenStudy (anonymous):

thanks, yeah using hopitals rule twice

OpenStudy (anonymous):

oops, the url broke

OpenStudy (anonymous):

that wolframalpha thing is realllly nice

OpenStudy (anonymous):

Oh damn, you're right. Man I suck at math.

OpenStudy (anonymous):

no, indeterminants are just really weird

OpenStudy (anonymous):

Harwin -- easy mistake to make. My students do it all the time.

OpenStudy (anonymous):

I was joking.

OpenStudy (anonymous):

wolframalpha.com (in case you haven't heard of it already)--it helps a lot with understanding IMO

OpenStudy (anonymous):

k thanks Esperantist, i'll check it out

OpenStudy (anonymous):

If anyone's still around I have sin(x)/[1-cos(x)], I got -1/2 again, is that right?

OpenStudy (anonymous):

what does it approach to?